OpenStudy (carlyleukhardt):

What is the solution to the equation 2k-10/-7 =8 ? k=?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\frac{2k-10}{-7}=8\]?

OpenStudy (carlyleukhardt):

hai :3

OpenStudy (carlyleukhardt):

Yes

OpenStudy (misty1212):

ok start with \[2k-10=-56\]

OpenStudy (misty1212):

add \(10\) then divide by \(2\)

OpenStudy (misty1212):

you got that?

OpenStudy (carlyleukhardt):

Not really

OpenStudy (misty1212):

ok lets to step step step

OpenStudy (carlyleukhardt):

okay ^_^

OpenStudy (misty1212):

\[\frac{2k-10}{-7}=8\]step one, multiply both sides by \(-7\) get \[2k-10=-56\]

OpenStudy (misty1212):

step two add \(10\) to both sides to get \[2j=-46\]

OpenStudy (misty1212):

oops \(2k=-46\)

OpenStudy (misty1212):

then divide both sides by 2 to get \[k=-23\]

OpenStudy (carlyleukhardt):

Thank you very very very much :D